# Story-Adapter: A Training-free Iterative Framework for Long Story Visualization

### v1

@@ Appendix B Subject-Consistent Generation Comparison
In the evaluation phase, we employ GPT-4o (OpenAI, 2024) according to the settings of StoryDiffusion (Zhou et al., 2024) to generate 20 character descriptions and 100 specific activity descriptions, respectively. We combine them as 2000 test descriptions, to compare Story-Adapter and subject-consistent image generation baselines, including IP-Adapter (Ye et al., 2023), PhotoMaker (Li et al., 2024), and StoryDiffusion (Zhou et al., 2024).
@@
Table 4: Quantitative comparison with subject-consistent image generation methods.
@@
Story-Adapter (Ours) 0.326 0.877
@@
Story-AdapterXL (Ours) 0.332 0.884
Quantitative Evaluation.
For quantitative comparisons on subject-consistent image generation, we employ CLIP text-to-image similarity (CLIP-T) and image-image similarity (CLIP-I) to measure consistency between the character images and generated images. Tab. 4 shows that Story-Adapter achieves SoTA performance in terms of both quantitative metrics, which demonstrates Story-Adapter’s ability to generate subject-consistent image sequences based on text prompts or image prompts.
Qualitative Evaluation.
Fig. 10 shows the qualitative comparison results. Story-Adapter generates higher-quality images in subject-consistency and detailed interactions. In contrast, IP-Adapter fails to generate correctly, e.g., “paper”, “whiteboard”, and “chainsaw”. PhotoMaker cannot generate images consistently, e.g., maintaining details of the attire. Despite accurately generating content according to text prompts with visual consistency, StoryDiffusion suffers from visualizing complex details due to lacking global story comprehension. By incorporating a global story view in our iterative paradigm, Story-Adapter can maintain visual consistency, especially in details throughout the story.
Appendix C Human Evaluation
Setting.
To complement the evaluation metrics to accurately reflect the quality of the generated stories, we involve human evaluation to further compare Story-Adapter and baselines. Referring to the setting in StoryGen (Liu et al., 2024), we invite participants to rate various aspects: text-image alignment (Align.), character interaction (Inter.), content consistency (Cons.), image quality (Qual.), and preference (Pref.) on a scale from 1 to 5. The higher the better.
Subject-Consistent Image Generation
Model Align. $\uparrow$ Inter. $\uparrow$ Cons. $\uparrow$ Qual. $\uparrow$ Pref. $\uparrow$
IP-Adapter (Ye et al., 2023) 2.51 3.27 4.58 4.33 4.19
IP-AdapterXL (Ye et al., 2023) 2.66 3.36 4.72 4.51 4.26
PhotoMaker (Li et al., 2024) 3.79 4.18 4.25 4.47 4.11
StoryDiffusion (Zhou et al., 2024) 4.15 4.28 4.50 4.54 4.48
Story-Adapter 4.02 4.20 4.41 4.39 4.33
Story-AdapterXL 4.20 4.35 4.58 4.61 4.54
Regular-Length Story Visualization
SDM (Rombach et al., 2022) 4.11 2.37 2.01 4.17 1.14
Prompt-SDM (Rombach et al., 2022) 4.03 3.49 1.99 4.40 1.26
Finetuned-SDM (Rombach et al., 2022) 3.35 3.82 2.15 3.41 1.60
AR-LDM (Pan et al., 2024) 3.08 3.64 2.90 2.64 2.05
StoryGen (Liu et al., 2024) 3.72 4.17 3.83 3.79 3.39
StoryDiffusion (Zhou et al., 2024) 3.96 4.48 4.52 4.24 4.37
Story-Adapter 3.89 4.21 4.36 4.09 4.10
Story-AdapterXL 4.06 4.60 4.74 4.53 4.62
Long Story Visualization
AR-LDM (Pan et al., 2024) 3.30 3.68 3.42 2.15 3.27
StoryGen (Liu et al., 2024) 3.51 4.06 3.88 2.72 3.51
IP-Adapter (Ye et al., 2023) 3.79 4.27 4.30 4.19 4.06
IP-AdapterXL (Ye et al., 2023) 3.83 4.23 4.61 4.47 4.11
StoryDiffusion (Zhou et al., 2024) 4.16 4.30 4.53 4.33 4.35
Story-Adapter 3.97 4.15 4.42 4.19 4.29
Story-AdapterXL 4.35 4.47 4.70 4.62 4.65
Results.
Tab. 5 shows that our Story-Adapter receives more preference from the participants. It is worth noting that although IP-Adapter receives higher scores for consistency in the subject-consistent image generation task, Story-Adapter is more favored in text-image alignment and generating character interactions. For regular-length and long story visualization, Story-Adapter is more preferred compared to baselines in most evaluation aspects, especially visual consistency and capability to generate character interactions. This is aligned with the quantitative measurement.
Appendix D More Iterations
Setting.
Figure 11: Story visualization results from different iterations by Story-Adapter. Accurate interactions are denoted in green, and wrong or missing ones are in red.
Results.
Tab. 6 shows that as iteration increases, Story-Adapter achieves significant improvement in visual consistency (aCCs and aFID) while text-image alignment (CLIP-T) drops slightly. This further demonstrates the contribution of the iterative paradigm to the semantic consistency of the overall story. However, we also note that a further increase in iterations harms text-image alignment, with limited gain in visual content consistency. This indicates that while Global Reference Cross-Attention (GRCA) effectively improves the content consistency of the long story, the increasing weighting factor of GRCA during the iterations poses a challenge to aligning the text prompts.
@@ Results.
Fig. 11 demonstrates a significant improvement in generative quality for fine-grained interactions as the iteration proceeds. The iterative paradigm effectively alleviates the diffusion model’s limitations on complex interaction generation by continuously creating input channels for text prompts. But more iterations wouldn’t improve the generation quality further. Therefore, 10 iterations in the iterative paradigm is an optimal choice based on the quantitative and qualitative experiments.
Appendix E More Visualization Results
In this section, we provide more visualization results from Story-Adapter and the baselines.
E.1 Visual Comparison
We compare the long story visualization results of representative work with AR-based, RI-based, and iterative paradigms, respectively. Specifically, Fig. 12, Fig. 13, and Fig. 14 show the generated results of the same “Pianist” story from the proposed Story-Adapter (iterative), StoryGen (Liu et al., 2024) (AR-based), and StoryDiffusion (Zhou et al., 2024) (RI-based), respectively.
@@
Fig. 12 shows that the visualization quality from StoryGen constantly gets worse as the length of the story increases. In Fig. 13, StoryDiffusion maintains high visual quality throughout the story, but it suffers from the flaw in the beginning frame that serves as the reference image, e.g., “closed-eye”. In addition, the subject “the character 1900” is not consistently generated as baby, kid, and adult. In contrast, our Story-Adapter effectively achieves high-quality story visualization and addresses the aforementioned limitations (see Fig. 14).
@@
Figure 12: Visualization results of StoryGen for the “Pianist” story. Zoom in for a better view.
Figure 13: Visualization results of StoryDiffusion for the “Pianist” story. Zoom in for a better view.
@@
Figure 14: Visualization results of our Story-Adapter for “Pianist”. Zoom in for a better view.
E.2 Longer Story Visualization Results
@@
In Fig. 15, we show the visualization results of the long story (up to 100 frames).
@@
E.3 Different Style
We provide the long story visualization results from Story-Adapter in a realistic style in Fig. 16. The experiment results suggest that Story-Adapter can be applied to different visual styles as well.

### v2

@@ Appendix E Subject-Consistent Generation Comparison
In the evaluation phase, we employ GPT-4o (OpenAI, 2024) according to the settings of StoryDiffusion (Zhou et al., 2024) to generate 20 character descriptions and 100 specific activity descriptions, respectively. We combine them as 2000 test descriptions, to compare Story-Iter and subject-consistent image generation baselines, including IP-Adapter (Ye et al., 2023), PhotoMaker (Li et al., 2024b), and StoryDiffusion (Zhou et al., 2024).
Quantitative Evaluation.
For quantitative comparisons on subject-consistent image generation, we employ CLIP text-to-image similarity (CLIP-T) and image-image similarity (CLIP-I) to measure consistency between the character images and generated images. Tab. 11 shows that Story-Iter achieves SoTA performance in terms of both quantitative metrics, which demonstrates Story-Iter’s ability to generate subject-consistent image sequences based on text prompts or image prompts.
Qualitative Evaluation.
Fig. 18 shows the qualitative comparison results. Story-Iter generates higher-quality images in subject-consistent and detailed interactions. In contrast, IP-Adapter fails to generate correctly, e.g., “paper”, “whiteboard”, and “chainsaw”. PhotoMaker cannot generate images consistently, e.g., maintaining details of the attire. Despite accurately generating content according to text prompts with visual consistency, StoryDiffusion suffers from visualizing complex details due to lacking global story comprehension. By incorporating a global story view in our iterative paradigm, Story-Iter can maintain visual consistency, especially in details throughout the story.
Figure 19: Story visualization results from different iterations by Story-Iter. Accurate interactions are denoted in green, wrong or missing ones are in red.
@@
Table 10: Quantitative comparison in subject-consistent image generation tasks.
@@
Story-Iter (Ours) 0.326 0.877
@@
Story-IterXL (Ours) 0.332 0.884
@@
Appendix F More Iterations
Setting.
@@ Results.
Tab. 11 shows that as iteration increases, Story-Iter achieves significant improvement in visual consistency (aCCs and aFID) while text-image alignment (CLIP-T) drops slightly. This further demonstrates the contribution of the iterative paradigm to the semantic consistency of the overall story. However, we also note that a further increase in iterations harms text-image alignment, with limited gain in visual content consistency. This indicates that while Global Reference Cross-Attention (GRCA) effectively improves the content consistency of the long story, the increasing weighting factor of GRCA during the iterations poses a challenge to aligning the text prompts.
Fig. 19 demonstrates a significant improvement in generative quality for fine-grained interactions as the iteration proceeds. The iterative paradigm effectively alleviates the diffusion model’s limitations on complex interaction generation by continuously creating input channels for text prompts. But more iterations wouldn’t improve the generation quality further. Therefore, based on the above results, fewer than 10 iterations are sufficient to achieve high-quality story visualization, with any further iterations yielding negligible improvements.
This trade-off between consistency and text-image alignment is analogous to autoregressive models, where increased sequence length may introduce compounding errors, yet also allows for stronger global coherence. This tradeoff does not indicate inefficiency in our iterative refinement process, but rather highlights the importance of balancing local fidelity and global consistency. Importantly, this trade-off can be effectively mitigated. Also, as demonstrated in Tab. 11, adjusting the hyperparameters based on the specific story length and content, as well as terminating the external iteration of Story-Iter earlier, can offer a better trade-off. This opens promising directions for future work, where we plan to integrate content-aware and length-sensitive hyperparameter scheduling, guided by automatic evaluations (e.g., GPT-based metrics), to further optimize both consistency and alignment.
Appendix G Limitation
While Story-Iter demonstrates great potential in long-story visualization, there remains an inherent trade-off between consistency and diversity. Although our initialization and linear weighting strategy help maintain global diversity throughout the story, certain local frames still exhibit diversity limitations. In future work, we aim to incorporate pose or layout conditional controls to further encourage diversity between consecutive frames.
At the same time, multiple external iterations for long story generation may affect the practical efficiency of Story-Iter. However, this limitation can be fundamentally addressed through accelerated generation methods such as MeanFlow (Geng et al., 2025), which we discuss in Sec. A.5.
Additionally, we observe a trade-off between iterations and text–image alignment: longer iterations enhance global consistency but weaken alignment. This does not indicate inefficiency of the iterative refinement process, but rather reflects the balance between local fidelity and global consistency. Notably, this trade-off can be mitigated by tuning hyperparameters or terminating iterations earlier, as discussed in Sec. F. For future work, we plan to introduce Qwen3-VL–based content-aware scheduling, where Qwen3-VL evaluates story semantics during refinement and adaptively decides: whether an iteration should stop early, and set the next iteration’s GRCA/text-conditioned weighting. This content-adaptive controller will further optimize consistency–alignment trade-offs without sacrificing Story-Iter’s strong narrative coherence. In addition, we plan to expand our benchmark and release a larger corpus to support more comprehensive and standardized evaluation for the community.
Appendix H More Visualization Results
In this section, we provide more visualization results from Story-Iter and the baselines.
H.1 Visual Comparison
We compare the long story visualization results of representative work with AR-based, RI-based, and iterative paradigms, respectively. Specifically, Fig. 24, Fig. 20, Fig. 21, Fig. 23, and Fig. 22 show the generated results of the same “Robinson” story from the proposed Story-Iter (iterative), StoryGen (Liu et al., 2024) (AR-based), IP-Adapter (Ye et al., 2023) (subject-consistent image generation), ConsiStory (Tewel et al., 2024) (RI-based), and StoryDiffusion (Zhou et al., 2024) (RI-based), respectively.
Results.
@@
Fig. 20 shows that the visualization quality from StoryGen constantly gets worse as the length of the story increases. The visualization results of the IP-Adapter in Fig. 21 show obvious interaction missing or wrong interactions. In Fig. 22, StoryDiffusion maintains high visual quality throughout the story but suffers from some shortcomings in multi-character interactions. In contrast, our Story-Iter effectively achieves high-quality story visualization and addresses the aforementioned limitations (see Fig. 24).
We also observe that, although Story-Iter employs global embeddings, it still preserves fine-grained consistency compared to methods relying heavily on local visual embeddings—such as StoryDiffusion, StoryGen, and ConsiStory. By summarizing an entire image into a compact representation, global embeddings enable computationally affordable processing while preserving subject relevance and semantic consistency across full-length story sequences. At the same time, our method leverages latent denoising features that retain pixel-wise, fine-grained visual information, guided by the global embedding. This combination helps preserve important local details, such as clothing and facial features, as illustrated in Fig. 24. In contrast, StoryDiffusion, StoryGen, and ConsiStory paradoxically yield inconsistent details, as they are substantially disrupted by irrelevant noise due to their excessive emphasis on local features.
@@
Figure 20: Visualization results of StoryGen for the “Robinson” story. Zoom in for a better view.
@@
Figure 21: Visualization results of IP-Adapter for the “Robinson” story. Zoom in for a better view.
Figure 22: Visualization results of StoryDiffusion for the “Robinson” story. Zoom in for a better view.
Figure 23: Visualization results of ConsiStory for the “Robinson” story. Zoom in for a better view.
Figure 24: Visualization results of our Story-Iter for “Robinson”. Zoom in for a better view.
H.2 Story-Iter-Fast
We provide qualitative long-story visualization result of Story-Iter-Fast in Fig. 25. This result demonstrates that the proposed GRCA mechanism and iterative framework are fully compatible with distilled and accelerated diffusion models; even under extremely efficient sampling regimes, Story-Iter-Fast is able to maintain strong global story coherence.
@@ H.2 Story-Iter-Fast
Figure 25: Long story visualization results of Story-Iter-Fast for “oyal dog”.
H.3 Different Style
We visualize scenes of the same long story in different styles in Fig. 26, Fig. 27, and Fig. 28. Meanwhile, we provide more long story visualization examples from Story-Iter in different styles in Fig. 29, Fig. 30, and Fig. 31. The experiment results suggest that Story-Iter can be applied to different visual styles as well. Furthermore, these results demonstrate that Story-Iter consistently generates high-quality visual narratives across a variety of initialization conditions—including different styles, content types, story lengths, and random seeds—highlighting its robustness and generalizability.
Figure 26: Our comic style story visualization results for “Robinson”. Zoom in for a better view.
Figure 27: Our realistic style story visualization results for “Robinson”. Zoom in for a better view.
Figure 28: Our film style story visualization results for “Robinson”. Zoom in for a better view.
Figure 30: Our monochrome style regular story visualization results in StorySalon (Liu et al., 2024). Zoom in for a better view.
Figure 31: Our realistic style story visualization results for “The Prince and the Princess”. Zoom in for a better view.
H.4 Longer Story Visualization Results
In Fig. 32 and Fig. 33, we show the visualization results of the long story (up to 100 frames).
Figure 32: Our long story visualization results for “Tortoise and the Hare race”. Zoom in for a better view.
@@ H.4 Longer Story Visualization Results
Appendix I Declaration of LLM Tool Usage
During the preparation of this manuscript, I used OpenAI’s GPT-4.1 model for minor language refinement and smoothing of the writing. The LLM tool was not used for generating original content, conducting data analysis, or formulating core scientific ideas. All conceptual development, experimentation, and interpretation were conducted independently without reliance on LLM tools. The other points involving the use of LLMs have already been highlighted in the paper.